# ProTeX: Structure-In-Context Reasoning and Editing of Proteins with Large Language Models

### v1

@@ Abstract
Large language models have made remarkable progress in the field of molecular science, particularly in understanding and generating functional small molecules. This success is largely attributed to the effectiveness of molecular tokenization strategies. In protein science, the amino acid sequence serves as the sole tokenizer for LLMs. However, many fundamental challenges in protein science are inherently structure-dependent. The absence of structure-aware tokens significantly limits the capabilities of LLMs for comprehensive biomolecular comprehension and multimodal generation. To address these challenges, we introduce a novel framework, ProTeX, which tokenizes the protein sequences, structures, and textual information into a unified discrete space. This innovative approach enables joint training of the LLM exclusively through the Next-Token Prediction paradigm, facilitating multimodal protein reasoning and generation. ProTeX enables general LLMs to perceive and process protein structures through sequential text input, leverage structural information as intermediate reasoning components, and generate or manipulate structures via sequential text output. Experiments demonstrate that our model achieves significant improvements in protein function prediction, outperforming the state-of-the-art domain expert model with a twofold increase in accuracy. Our framework enables high-quality conformational generation and customizable protein design. For the first time, we demonstrate that by adopting the standard training and inference pipelines from the LLM domain, ProTeX empowers decoder-only LLMs to effectively address diverse spectrum of protein-related tasks.
keywords:
American Chemical Society,
[…]
{tocentry}
Some journals require a graphical entry for the Table of Contents. This should be laid out “print ready” so that the sizing of the text is correct.
Inside the tocentry environment, the font used is Helvetica 8 pt, as required by Journal of the American Chemical Society.
The surrounding frame is 9 cm by 3.5 cm, which is the maximum permitted for Journal of the American Chemical Society graphical table of content entries. The box will not resize if the content is too big: instead it will overflow the edge of the box.
This box and the associated title will always be printed on a separate page at the end of the document.
@@ 1.   Introduction
Inspired by the multi-modal LLMs Emu3 29 and Chamelon 30, we present ProTeX, a dedicated framework for tokenizing both the 1D sequence and 3D structure of proteins for LLMs, akin to SMILES. Just as TeX provides precise control over document layout and formatting, ProTeX empowers researchers for protein data formatting and editing, which enables LLMs to reason and generate sequences that arbitrarily interleave textual and protein modality. ProTeX can employ the advanced Reasoning paradigm, CoT reasoning, to enhance LLMs’ capability in protein-related tasks. This approach allows structural or textual information to serve as a key logical component in the reasoning process, making protein problem solving more transparent, logical, interpretable, and controllable. We have constructed a unified foundational framework that leverages a single model and loss function based solely on the next-token prediction (NTP) strategy, enabling seamless adaptation to diverse protein-related downstream tasks. In comparison to the state-of-the-art domain expert LLMs BiomedGPT 23 and Llama2-molinst-protein-7B 31, our approach demonstrates a twofold enhancement in accuracy for protein function prediction. Our findings reveal that the CoT approach and sampling strategies from the field of LLM can be effortlessly adapted to the protein domain, enabling high-accuracy protein structure prediction and conformational sampling. Additionally, the integration of arbitrary textual inputs allows for controllable protein generation based on human-defined prompts. For the first time, we demonstrate that decoder-only LLMs are capable of understanding, predicting, and designing proteins.
@@ 2.   Methods
ProTeX represents protein sequences, structures, and natural language text as a series of reversible discrete tokens, leveraging the unified training paradigm of auto-regressive transformers32, 33. During training, we construct various prompts with arbitrary orders of protein and text, enabling the model to complete different downstream tasks, ranging from unimodal inference to multimodal CoT generation. The main architechture of the model is shown in Figure  1A.
@@ 2.1.   Tokenizing All-Atom Protein Structures
ProTeX tokenizes the protein backbone structure following the work of Lin et al. 34. Here, we outline the main technical strategy. The orginal work focuses on tokenizing the metastable conformational structure of $x$ of proteins into discrete tokens through a model that includes three main components: the encoder $f_{\theta}$, the tokenizer $h_{\theta}$, and the decoder $g_{\phi}$. The original training objective is simply to reconstruct the encoded structures.
@@ 2.3.1 Dataset
We first curated a database of 3.36 million proteins, including their sequences and structures, from the clustered AlphaFold Protein Structure Database (AFDB) v4 dataset 39, the Swiss-Prot database released in May 2022 40, and RCSB PDB 41. All of the proteins were released before July 25, 2022. We then processed this database using ProTeX introduced in Section 2.1, for structural reconstruction and filtering to obtain the sequence and structural tokens for every protein. The dataset is subsequently split into training (90%), validation (5%), and test (5%) sets.
@@ 2.3.4 Baseline Setups
  • $\text{ProTeX}_{\text{Llama3}}$ Our main proposed model, which is supervised fine-tuned on Meta-Llama-3-8B-Base using the ProTeX tokenizer with our total dataset, which incorporates training data from different downstream tasks. The textual protein sequence information in PFUD dataset is replaced by our tokenized protein sequence and structure. We also conducted ablation experiments under different training conditions, including various training strategies such as Low-Rank Adaptation (LoRA) 49, different training sets, and model scales. For more details on the ablation experiments, please refer to Table 4, Table 3, and Section A.4.
@@ 3.   Results and discussion
3.1.   ProTeX Enables Structure-In-Context Protein Understanding
@@
As shown in Table 2, our analysis demonstrates that the multi-task fine-tuned model $\text{ProTeX}_{\text{Llama3}}$ achieves optimal performance, excelling in both linguistic fluency and accuracy in addressing domain-specific professional protein queries, highlighting the effectiveness of our ProTeX framework and the existence of overlapping subspaces among different tasks which facilitate mutual enhancement during the training process. Furthermore, we conduct a comparative analysis by fine-tuning models exclusively on the PFUD dataset to evaluate their respective performance in the ablation study A.4. The experimental results in Table 2 and Table 4 demonstrate that incorporating the novel ProTeX tokenizer significantly outperforms the the fine-tuned models that use one-letter amino acid abbreviations (Llama2-molinst-protein-7B,Llama3-AAseq-FT) or compositional protein encoders (ProtT3-FT, BiomedGPT). These results highlights the limitations of current models in functional understanding when relying solely on textual sequence alphabets or other compositional approaches, further substantiating the beneficial impact of incorporating structural tokens in enhancing protein function comprehension. On our PFUD test set, where all QA pairs are derived exclusively from Swiss-Prot proteins, the ROUGE-L score of the Llama2-molinst-protein-7B 31 model closely matches its officially reported metrics. However, we find that ROUGE-L is not a robust indicator for assessing the precision of key terms, as demonstrated by the substantially lower exact match performance. Our ablation studies presented in Table 4 demonstrate that pre-training without explicit functional information significantly enhances functional comprehension. These results collectively suggest a symbiotic relationship between multimodal understanding and generation. Specifically, enhancing comprehension data improves generative task performance, and conversely, expanding generative training data strengthens interpretative capabilities. This phenomenon aligns with recent observations in computer vision research 54. Additionally, we evaluated domain-specific response performance, as illustrated in Figure 2. Although the fine-tuned ProtT3 model achieves comparable performance with our proposed model in motif recognition and multi-attribute tasks, ProTeX consistently outperforms competing approaches across various domains. Notably, the BiomedGPT model exhibits suboptimal performance throughout our experiments. Despite extensive efforts to optimize the prompt formulation by adopting BiomedGPT’s preferred structure, the model demonstrates substantial limitations in addressing a wide range of protein-related queries. This suggests that BiomedGPT’s training may be insufficient for comprehensive protein-related tasks. Moreover, approximately 2% of BiomedGPT’s responses yielded unknown answers, indicating that the external encoder integration strategy may cause substantial representational shifts in response to sequence similarity variations, potentially increasing perplexity and reducing overall performance. These observations underscore the effectiveness of our early fusion training strategy 55, 56, 57, which integrates interleaved textual and protein modalities into a unified representation, enabling natural and intrinsic connections between different modalities.
@@
$\text{ProTeX}_{\text{Llama3}}$ 71.73 41.54 63.46 43.17 57.89
@@
$\text{ProTeX}_{\text{Llama3}}$ (w/o Multi-Task) 1B ProTeX 1B - PFUD Full
$\text{ProTeX}_{\text{Llama3}}$ (w/o Multi-Task) lora ProTeX 8B - PFUD Lora
$\text{ProTeX}_{\text{Llama3}}$ (w/o Multi-Task) ProTeX 8B - PFUD Full
$\text{ProTeX}_{\text{Llama3}}$ (w/o Multi-Task) w/ PT ProTeX 8B PSPD PFUD Full
$\text{ProTeX}_{\text{Llama3}}$ (Proposed) ProTeX 8B - PFUD,PSPD, PDD, PSAD Full
@@
$\text{ProTeX}_{\text{Llama3}}$ (w/o Multi-Task) 1B 64.97 39.40 61.25 40.56 55.09
$\text{ProTeX}_{\text{Llama3}}$ (w/o Multi-Task) lora 62.13 38.80 60.38 39.60 53.72
$\text{ProTeX}_{\text{Llama3}}$ (w/o Multi-Task) 66.12 40.01 62.87 41.71 56.15
$\text{ProTeX}_{\text{Llama3}}$ (w/o Multi-Task) w/ PT 70.57 40.39 63.77 42.69 57.17
$\text{ProTeX}_{\text{Llama3}}$ (Proposed) 71.73 41.54 63.46 43.17 57.89
@@
3.2.   ProTeX Enables Structure-Involved Reasoning for Proteins
@@
Figure 3: Multimodal chain-of-thought with multi-round chat. (A) Direct Prompting, direct asking the question of protein structure or protein function. (B) Chain-of-Thought Prompting, first analyzes the sequence, then generates the structure, and subsequently infers the function step-by-step.
@@ 3.4.   Knowledge-guided and natural language-instructed protein design
Figure 6: (A) Design chat and results for (A) dITP/XTP pyrophosphatase and (B) cytidylate kinase.
@@ 4.   Concluding Remarks
In this paper, we present a novel unified framework which achieves dual innovation in protein science. (1) Architectural unification of core challenges through a foundation model with task-agnostic formulation via a single-model, unified-loss paradigm. (2) Pioneering the application of Chain-of-Thought reasoning in the multimodal protein reasoning and generation. ProTeX enables LLMs to effectively process protein-related tasks through a mixed-modal fusion strategy. By employing in-context supervised fine-tuning, we have successfully integrate domain-specific knowledge into general-purpose LLMs, such as Llama3, equipping them with preliminary capabilities in multimodal protein comprehension and generation. Our model not only demonstrates the ability to tackle a wide range of protein tasks, including protein understanding, structure generation and protein design, but also introduces multimodal Chain-of-Thought reasoning, enhancing the transparency of the model’s deductive processes and ensuring greater interpretability.
Our current model may exhibit slight performance gaps in certain tasks compared to task-specific or domain-expert models, such as ESMFold2, 1. Notably, ESMFold’s success is largely attributed to its massive parameter count and extensive training dataset, which surpass even those of Llama3 8B in scale. Considering the well-known scaling laws of LLMs, we can anticipate continuous performance improvements across various tasks with the development of larger LLMs and more biological data. Beyond scaling parameters, well-established paradigms in the LLM domain, such as reinforcement learning-based alignment65 and inference-time self-improvement66, can be systematically applied to enhance ProTeX’s performance across various protein-related tasks. Exploring these approaches will be a key focus for our future research.
@@ A.1.   Data and Software Availability
Our model and dataset will be publicly available on Hugging Face:https://huggingface.co/mzcwd/ProTeX
@@ A.2.   Dataset
We then process this dataset using the ProTeX tokenizer, introduced in Section 2.1, for structural reconstruction, obtaining a tokenized string for each protein. Proteins with a reconstructed TM-score above 90 are retained. The dataset is subsequently split into training (90%), validation (5%), and test (5%) sets. For all protein-related questions constructed in the following sections, protein accessions are sourced from UniProt. Accordingly, problem-specific datasets are created by matching the corresponding accessions from the training, validation, and test sets. If an accession is not found in our database, the corresponding question is excluded from the dataset.

### v2

# ProtTeX: Structure-In-Context Reasoning and Editing of Proteins with Large Language Models

@@ Abstract
Large language models have made remarkable progress in the field of molecular science, particularly in understanding and generating functional small molecules. This success is largely attributed to the effectiveness of molecular tokenization strategies. In protein science, the amino acid sequence serves as the sole tokenizer for LLMs. However, many fundamental challenges in protein science are inherently structure-dependent. The absence of structure-aware tokens significantly limits the capabilities of LLMs for comprehensive biomolecular comprehension and multimodal generation. To address these challenges, we introduce a novel framework, ProtTeX, which tokenizes the protein sequences, structures, and textual information into a unified discrete space. This innovative approach enables joint training of the LLM exclusively through the Next-Token Prediction paradigm, facilitating multimodal protein reasoning and generation. ProtTeX enables general LLMs to perceive and process protein structures through sequential text input, leverage structural information as intermediate reasoning components, and generate or manipulate structures via sequential text output. Experiments demonstrate that our model achieves significant improvements in protein function prediction, outperforming the state-of-the-art domain expert model with a twofold increase in accuracy. Our framework enables high-quality conformational generation and customizable protein design. For the first time, we demonstrate that by adopting the standard training and inference pipelines from the LLM domain, ProtTeX empowers decoder-only LLMs to effectively address diverse spectrum of protein-related tasks.
@@ 1.   Introduction
Inspired by the multi-modal LLMs Emu3 29 and Chamelon 30, we present ProtTeX, a dedicated framework for tokenizing both the 1D sequence and 3D structure of proteins for LLMs, akin to SMILES. Just as TeX provides precise control over document layout and formatting, ProtTeX empowers researchers for protein data formatting and editing, which enables LLMs to reason and generate sequences that arbitrarily interleave textual and protein modality. ProtTeX can employ the advanced Reasoning paradigm, CoT reasoning, to enhance LLMs’ capability in protein-related tasks. This approach allows structural or textual information to serve as a key logical component in the reasoning process, making protein problem solving more transparent, logical, interpretable, and controllable. We have constructed a unified foundational framework that leverages a single model and loss function based solely on the next-token prediction (NTP) strategy, enabling seamless adaptation to diverse protein-related downstream tasks. In comparison to the state-of-the-art domain expert LLMs BiomedGPT 23 and Llama2-molinst-protein-7B 31, our approach demonstrates a twofold enhancement in accuracy for protein function prediction. Our findings reveal that the CoT approach and sampling strategies from the field of LLM can be effortlessly adapted to the protein domain, enabling high-accuracy protein structure prediction and conformational sampling. Additionally, the integration of arbitrary textual inputs allows for controllable protein generation based on human-defined prompts. For the first time, we demonstrate that decoder-only LLMs are capable of understanding, predicting, and designing proteins.
@@ 2.   Methods
ProtTeX represents protein sequences, structures, and natural language text as a series of reversible discrete tokens, leveraging the unified training paradigm of auto-regressive transformers32, 33. During training, we construct various prompts with arbitrary orders of protein and text, enabling the model to complete different downstream tasks, ranging from unimodal inference to multimodal CoT generation. The main architechture of the model is shown in Figure  1A.
@@ 2.1.   Tokenizing All-Atom Protein Structures
ProtTeX tokenizes the protein backbone structure following the work of Lin et al. 34. Here, we outline the main technical strategy. The orginal work focuses on tokenizing the metastable conformational structure of $x$ of proteins into discrete tokens through a model that includes three main components: the encoder $f_{\theta}$, the tokenizer $h_{\theta}$, and the decoder $g_{\phi}$. The original training objective is simply to reconstruct the encoded structures.
@@ 2.3.1 Dataset
We first curated a database of 3.36 million proteins, including their sequences and structures, from the clustered AlphaFold Protein Structure Database (AFDB) v4 dataset 39, the Swiss-Prot database released in May 2022 40, and RCSB PDB 41. All of the proteins were released before July 25, 2022. We then processed this database using ProtTeX introduced in Section 2.1, for structural reconstruction and filtering to obtain the sequence and structural tokens for every protein. The dataset is subsequently split into training (90%), validation (5%), and test (5%) sets.
@@ 2.3.4 Baseline Setups
  • $\text{ProtTeX}_{\text{Llama3}}$ Our main proposed model, which is supervised fine-tuned on Meta-Llama-3-8B-Base using the ProtTeX tokenizer with our total dataset, which incorporates training data from different downstream tasks. The textual protein sequence information in PFUD dataset is replaced by our tokenized protein sequence and structure. We also conducted ablation experiments under different training conditions, including various training strategies such as Low-Rank Adaptation (LoRA) 49, different training sets, and model scales. For more details on the ablation experiments, please refer to Table 4, Table 3, and Section A.4.
@@ 3.   Results and discussion
3.1.   ProtTeX Enables Structure-In-Context Protein Understanding
@@
As shown in Table 2, our analysis demonstrates that the multi-task fine-tuned model $\text{ProtTeX}_{\text{Llama3}}$ achieves optimal performance, excelling in both linguistic fluency and accuracy in addressing domain-specific professional protein queries, highlighting the effectiveness of our ProtTeX framework and the existence of overlapping subspaces among different tasks which facilitate mutual enhancement during the training process. Furthermore, we conduct a comparative analysis by fine-tuning models exclusively on the PFUD dataset to evaluate their respective performance in the ablation study A.4. The experimental results in Table 2 and Table 4 demonstrate that incorporating the novel ProtTeX tokenizer significantly outperforms the the fine-tuned models that use one-letter amino acid abbreviations (Llama2-molinst-protein-7B,Llama3-AAseq-FT) or compositional protein encoders (ProtT3-FT, BiomedGPT). These results highlights the limitations of current models in functional understanding when relying solely on textual sequence alphabets or other compositional approaches, further substantiating the beneficial impact of incorporating structural tokens in enhancing protein function comprehension. On our PFUD test set, where all QA pairs are derived exclusively from Swiss-Prot proteins, the ROUGE-L score of the Llama2-molinst-protein-7B 31 model closely matches its officially reported metrics. However, we find that ROUGE-L is not a robust indicator for assessing the precision of key terms, as demonstrated by the substantially lower exact match performance. Our ablation studies presented in Table 4 demonstrate that pre-training without explicit functional information significantly enhances functional comprehension. These results collectively suggest a symbiotic relationship between multimodal understanding and generation. Specifically, enhancing comprehension data improves generative task performance, and conversely, expanding generative training data strengthens interpretative capabilities. This phenomenon aligns with recent observations in computer vision research 54. Additionally, we evaluated domain-specific response performance, as illustrated in Figure 2. Although the fine-tuned ProtT3 model achieves comparable performance with our proposed model in motif recognition and multi-attribute tasks, ProtTeX consistently outperforms competing approaches across various domains. Notably, the BiomedGPT model exhibits suboptimal performance throughout our experiments. Despite extensive efforts to optimize the prompt formulation by adopting BiomedGPT’s preferred structure, the model demonstrates substantial limitations in addressing a wide range of protein-related queries. This suggests that BiomedGPT’s training may be insufficient for comprehensive protein-related tasks. Moreover, approximately 2% of BiomedGPT’s responses yielded unknown answers, indicating that the external encoder integration strategy may cause substantial representational shifts in response to sequence similarity variations, potentially increasing perplexity and reducing overall performance. These observations underscore the effectiveness of our early fusion training strategy 55, 56, 57, which integrates interleaved textual and protein modalities into a unified representation, enabling natural and intrinsic connections between different modalities.
@@
$\text{ProtTeX}_{\text{Llama3}}$ 71.73 41.54 63.46 43.17 57.89
@@
$\text{ProtTeX}_{\text{Llama3}}$ (w/o Multi-Task) 1B ProtTeX 1B - PFUD Full
$\text{ProtTeX}_{\text{Llama3}}$ (w/o Multi-Task) lora ProtTeX 8B - PFUD Lora
$\text{ProtTeX}_{\text{Llama3}}$ (w/o Multi-Task) ProtTeX 8B - PFUD Full
$\text{ProtTeX}_{\text{Llama3}}$ (w/o Multi-Task) w/ PT ProtTeX 8B PSPD PFUD Full
$\text{ProtTeX}_{\text{Llama3}}$ (Proposed) ProtTeX 8B - PFUD,PSPD, PDD, PSAD Full
@@
$\text{ProtTeX}_{\text{Llama3}}$ (w/o Multi-Task) 1B 64.97 39.40 61.25 40.56 55.09
$\text{ProtTeX}_{\text{Llama3}}$ (w/o Multi-Task) lora 62.13 38.80 60.38 39.60 53.72
$\text{ProtTeX}_{\text{Llama3}}$ (w/o Multi-Task) 66.12 40.01 62.87 41.71 56.15
$\text{ProtTeX}_{\text{Llama3}}$ (w/o Multi-Task) w/ PT 70.57 40.39 63.77 42.69 57.17
$\text{ProtTeX}_{\text{Llama3}}$ (Proposed) 71.73 41.54 63.46 43.17 57.89
@@
3.2.   ProtTeX Enables Structure-Involved Reasoning for Proteins
@@
Figure 3: Multimodal chain-of-thought with multi-round chat. (A) Direct Prompting, direct asking the question of protein structure or protein function. (B) Chain-of-Thought Prompting, first analyzes the sequence, then generates the structure, and subsequently infers the function step-by-step. The Llama icon is sourced from https://github.com/alexrozanski/LlamaChat.
@@ 3.4.   Knowledge-guided and natural language-instructed protein design
Figure 6: Design chat and results for (A) dITP/XTP pyrophosphatase and (B) cytidylate kinase.
@@ 4.   Concluding Remarks
In this paper, we present a novel unified framework which achieves dual innovation in protein science. (1) Architectural unification of core challenges through a foundation model with task-agnostic formulation via a single-model, unified-loss paradigm. (2) Pioneering the application of Chain-of-Thought reasoning in the multimodal protein reasoning and generation. ProtTeX enables LLMs to effectively process protein-related tasks through a mixed-modal fusion strategy. By employing in-context supervised fine-tuning, we have successfully integrate domain-specific knowledge into general-purpose LLMs, such as Llama3, equipping them with preliminary capabilities in multimodal protein comprehension and generation. Our model not only demonstrates the ability to tackle a wide range of protein tasks, including protein understanding, structure generation and protein design, but also introduces multimodal Chain-of-Thought reasoning, enhancing the transparency of the model’s deductive processes and ensuring greater interpretability.
Our current model may exhibit slight performance gaps in certain tasks compared to task-specific or domain-expert models, such as ESMFold2, 1. Notably, ESMFold’s success is largely attributed to its massive parameter count and extensive training dataset, which surpass even those of Llama3 8B in scale. Considering the well-known scaling laws of LLMs, we can anticipate continuous performance improvements across various tasks with the development of larger LLMs and more biological data. Beyond scaling parameters, well-established paradigms in the LLM domain, such as reinforcement learning-based alignment65 and inference-time self-improvement66, can be systematically applied to enhance ProtTeX’s performance across various protein-related tasks. Exploring these approaches will be a key focus for our future research.
@@ A.1.   Data and Software Availability
Our model and dataset will be publicly available on Hugging Face:https://huggingface.co/mzcwd/ProtTeX
@@ A.2.   Dataset
We then process this dataset using the ProtTeX tokenizer, introduced in Section 2.1, for structural reconstruction, obtaining a tokenized string for each protein. Proteins with a reconstructed TM-score above 90 are retained. The dataset is subsequently split into training (90%), validation (5%), and test (5%) sets. For all protein-related questions constructed in the following sections, protein accessions are sourced from UniProt. Accordingly, problem-specific datasets are created by matching the corresponding accessions from the training, validation, and test sets. If an accession is not found in our database, the corresponding question is excluded from the dataset.